OpenStudy (anonymous):

Solve: 0 = 4x2 − 8x − 5.

OpenStudy (saifoo.khan):

Try using the quadratic formula..

OpenStudy (zzr0ck3r):

i always try the AC method first, here we have AC = 4(-5) = -20 so we want to numbers that multiply to -20 but add to -8 (the b value). two such numbers are -10 and 2, then i expant the middle term using these numbers 4x^2 + 2x - 10x - 5 = 0 (notice is still the same thing that we started with) then I factor something(the most I can) out of the first two terms so 2x devides both 4x^2 and 2x 2x(2x+1) then I want to factor something out of the 3rd and 4th terms such that what I have left is the same as (2x+1) so 2x(2x+1) -2(2x+1) = 0 then couple the terms (2x+1)(2x-1) = 0 the solve 2x+1 = 0 x = -1/2 2x-1 = 0 x = 1/2

OpenStudy (anonymous):

\[+-b \sqrt{b2-4*a*c}/2*a\]

OpenStudy (zzr0ck3r):

All normal forms are good and work fine, the reason I tell people that are new to this to use the AC method is because it really lets you get a feel for factoring...which comes in handy later. I guess it depends on what you are doing in class.

OpenStudy (anonymous):

I like that way, but what do you do next to get the fina answer? because 1/2 & -1/2 aren'y an answer choice

OpenStudy (anonymous):

final*